find dipole and quadruple moments of a cube of side a carrying a total charge Q which is uniformly distributed for a cube centered at the origin


1
Expert's answer
2021-11-07T19:59:38-0500

Dipole moment

Electric field (E)="\frac{kq}{r^3}"

Potential (V)="\frac{kQ}{r^2}"



Quaderpole moment

Electric field

"E=\frac{kq}{r^4}"

Potential

"V=\frac{kQ}{r^3}"